Case I: Elasticity Part I Demand and Supply are represented by the functions below: QD = 8250 – 325P QS = 850 + 175P Exercise: 1. Compute quantity and price in equilibrium 2. Determine what would happen if the price changed to 12$ 3. Plot the graph representing the first two questions. 4. Compute the elasticity of both curves assuming the price would increase from 22$ to 24$. 5. Plot the graph for question 4.


1
Expert's answer
2021-11-02T17:50:09-0400

1.

At market equilibrium, 

   "Q\n_\nD\n\n=Q\n_\nS\n\n=Q^*"

 Putting values

    8250−325P=850+175P

⇒500P=7400

P*=14.8

Equilibrium quantity:

Q*=8250−325(14.8)=3440


2.

If price changed to $12

In this case the quantity demanded is greater than the quantity supply. 

At P=12

"Q\n\n_D\n\n=8250\u2212325(12) =4350"

and quantity supply

"Q\n\n_S\n\n=850+175(12)=2950"


3.


4

When price is 22

Quantity demanded 

"Q\n\nD\n\n=8250\u2212325(22) =1100"


When P =24

Qd=8250−325(24)=450

Price elasticity of demand is calculated as

"E=\frac{\%change\space in\space demand}{

\% change\space in\space price}"


"⇒E=\frac{\frac{Q_

2

−Q
_
1}{

Q

_1}
}{\frac{


P

_2

−P

_1}{

P
_
1}}"


"⇒E=\frac{\frac{450

−1100}{

1100}
}{\frac{


24

−22}{

22}}"


"⇒E=−6.5\\

⇒|


E|



=6.5"


So price elasticity of demand when price increase from 22 to 24 is 6.5


Elasticity of supply:

At price P=22, 

Qs= 850+175(22)=4700

When price is P=24

Qs=850+175*24=5050

Price elasticity of supply:

"E'=\frac{\frac{5050-4700}{4700}}{\frac{24-22}{22}}"

E'=0.819


5.